OpenStudy (anonymous):

solve please explain steps 2/3x + 5 = 9

OpenStudy (anonymous):

2/3x+5=9 Get all like terms onto one side So Minus 5 from each side 2/3x=4 Now divide by 2/3 Remember when you divide by fractions, use their reciprocals and multipy So it'll be 4/ (2/3)= 4 * (3/2) = 6

OpenStudy (anonymous):

subtract 5 from both sides then divide by 2/3 on both sides and you get your answer :)